A student takes out two loans, one at 5% interest and the other at 6% interest. The 5% loan is $1000 more than the 6% loan,and the total interest for one year is $215. How much is each loan?
Answer by robertb(5830)

0.05*(l+1000) + 0.06*l = 215
==> 0.05l + 50 + 0.06l = 215
==> 0.11l = 165
==> l = 1,500 $, the loan in the 6%, and
l + 1,000 = 2,500 $. the loan in the 5%.